OpenStudy (anonymous):

Solve the following system of equations, and round you answer to the nearest tenths. 5x+2y=7 -2x+6y=9

OpenStudy (anonymous):

y=59/34

OpenStudy (texaschic101):

5x + 2y = 7 --->(-3) -2x + 6y = 9 --------------- -15x - 6y = -21 (result of multiplying by -3) -2x + 6y = 9 --------------add -17x = - 12 x = 12/17 -2x + 6y = 9 -2(12/17) + 6y = 9 -24/17 + 6y = 9 6y = 9 + 24/17 6y = 153/17 + 24/17 6y = 177/17 y = (177/17) / 6 y = 177/17 * 1/6 y = 177/102 = 59/34 check.. -2x + 6y = 9 -2(12/17) + 6(59/34) = 9 -24/17 + 354/34 = 9 -48/34 + 354/34 = 9 306/34 = 9 9 = 9 (correct) x = 12/17 y = 59/34

OpenStudy (anonymous):

whoah, it's a lot of fractions, but it makes sense, thanks!:)